Radius =2.7 cm , Q=6.3 × 10^-8 and volume charge density = ?
1
Expert's answer
2020-03-23T12:22:55-0400
"\rho_C=\frac{Q}{V}"

"\rho_C=\frac{3Q}{4\pi R^3}"

"\rho_C=\frac{3(6.3\cdot 10^{-8})}{4\pi (0.027)^3}=7,6\cdot 10^{-4}\frac{C}{m^3}"